OpenStudy (anonymous):

What is the solution set of |2x + 4| < 12? What is the solution set of |4x – 3| – 1 > 12? What is the solution set of |x – 6| = 4?

OpenStudy (anonymous):

|f(x)| < p => -p < f(x) < p

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

Okay, let's start with: |x – 6| = 4 What are the two possible scenarios that you're going to have to solve to find x?

OpenStudy (anonymous):

add 6 on both sides?

OpenStudy (anonymous):

No, the absolute value is already isolated.

OpenStudy (anonymous):

|f(x)| = p => f(x) = p , f(x) = -p

OpenStudy (anonymous):

i honestly dont understand that @Coolsector :/

OpenStudy (anonymous):

whats F=?

OpenStudy (anonymous):

|x – 6| = 4 let x-6 = f(x) 4 = p so |f(x)| = p mean two options f(x) = p , f(x) = -p

OpenStudy (anonymous):

it think you get x-6 = -4 x-6 = 4 always remember that in sums having mode there is never one answer you get two or range of answers

OpenStudy (anonymous):

You'll want to remove the absolute value bars by writing two equations: x-6 = 4 (this one you get just by dropping the bars) x-6 = -4 (and the negative case.)

OpenStudy (anonymous):

Ok so whenever the absolute value is already isolated , you make an equation out of them to get your anwsers?

OpenStudy (anonymous):

yes............now try to solve first one

OpenStudy (anonymous):

Yeah, just drop the bars for the first case, and then for the second case, switch the signs if you're dealing with an inequality (you're not here) and make it negative.

OpenStudy (anonymous):

okay, so would x=10?

OpenStudy (anonymous):

Yes, so now for the second one.

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

x-6 = -4 +6 +6 x = 2

OpenStudy (anonymous):

okay i see how you got that where the <12 come into play?

OpenStudy (anonymous):

when |f(x)|<12 then -12<f(x)<12 try to use this

OpenStudy (anonymous):

Okay, now let's go through the first problem: |2x + 4| < 12 This one is slightly different. When you create your first of the two scenarios, you're going to just drop the absolute value bars. So, what's the first scenario?

OpenStudy (anonymous):

2x=4<12

OpenStudy (anonymous):

2x + 4 < 12 For the second scenario, you're going to switch the inequality sign AND make the 12 negative. So what's the second scenario?

OpenStudy (anonymous):

2x=4<-12

OpenStudy (anonymous):

You have to switch to inequality sign, so what once what < is now >.

OpenStudy (anonymous):

so it would be" 2x=4>-12

OpenStudy (anonymous):

-12<2x+4<12 sub -4 from both sides -16<2x<8 divide both sides with 2 -8<x<4

OpenStudy (anonymous):

Yep, but you've got an = where a + should be. Now solve both scenarios.

OpenStudy (anonymous):

@melbel is it correct?

OpenStudy (anonymous):

opps lol my bad! 2x+4=>-12

OpenStudy (anonymous):

Yeah, uzumakhi, you got it. So what did you get, @CIERA1338, when you solved each scenario. (I keep calling them scenarios. What are they really called @uzumakhi?)

OpenStudy (anonymous):

i get how @uzumakhi worked it previously but y does he flipped the equation?

OpenStudy (anonymous):

okay just noticed he didnt flip it lol he subtracted the 12< from both sides right to get rid of it?

OpenStudy (anonymous):

i dont know what they really call it but i know how to use it

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

-8<x<4 is what i would come out with in the end after dividing 2 into both sides

OpenStudy (anonymous):

check your answer..........i also got the same

OpenStudy (anonymous):

That's what I got.

OpenStudy (anonymous):

so only second part is left

OpenStudy (anonymous):

So what is the first step in solving this one: |4x – 3| – 1 > 12

OpenStudy (anonymous):

you gotta get the |4x-3| by itself right?

OpenStudy (anonymous):

|4x-3| -1 > 12 |4x-3| > 11 so we have two conditions |4x-3| > 11 |4x-3| > -11

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

|4x-3>11 |4x-3|>-11

OpenStudy (anonymous):

these are the 2 scenarios right? @melbel

OpenStudy (anonymous):

4x>14 x>14/4

OpenStudy (anonymous):

1. |4x-3|>11 2 . |4x-3|>-11 Yep, but remember in the second scenario (the negative case) to flip the inequality sign when you drop the absolute value bars. So drop ur bars. :P

OpenStudy (anonymous):

1.) 4x-3>11 2.) 4x-3<-11

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

yay! now i subtract 3 from both sides and 4x>8 for the first on right??

OpenStudy (anonymous):

Well, this time you'll add three to both sides (since it's negative)

OpenStudy (anonymous):

1.) 4x-3>11 +3 +3 (so that you can knock out that -3 on that one side) So you'll get 4x>14 for the first one (so far, anyway.)

OpenStudy (anonymous):

4x>14 4x<-14

OpenStudy (anonymous):

yes, i got the first one would the second one be the same or no?

OpenStudy (anonymous):

Careful on the second one: 2.) 4x-3<-11 You're adding a positive 3 to a -11 So it'll be -8 4x<-8 I do the same thing all the time, lol!

OpenStudy (anonymous):

opps!

OpenStudy (anonymous):

1.) 4x>14 2.) 4x<-8 Now we have to get x by itself for each of the scenarios.

OpenStudy (anonymous):

the first one kind of sucks, you have (4x>14)/4 x>14/4 simplify the fraction x>7/2 (I hate fractions.)

OpenStudy (anonymous):

so you'd get like 2.5

OpenStudy (anonymous):

The second one is nicer: 4x<-8 x<-2

OpenStudy (anonymous):

3.5 and it depends whether or not your teacher allows decimals. Mine punishes us with fractions, so I would have to leave it at x>7/2

OpenStudy (anonymous):

x>7/2 x>3.5 not 2.5

OpenStudy (anonymous):

i was just going by my anwser chioces lol x > –2.5 or x > 4 x < –2.5 or x > 4 x < 2.5 or x > 4 x < –4 or x > 4

OpenStudy (anonymous):

what would i get as the anwser?